OpenStudy (anonymous):

Solve 2x/x-5 = 1/x+6

OpenStudy (solomonzelman):

if x-5 and x+6 are the denominators... CROSS MULTIPLY and SOLVE FOR X.

OpenStudy (anonymous):

this is a small world XDD

OpenStudy (anonymous):

Haha yeah can you help me

OpenStudy (anonymous):

I tried cross multiplying but didnt get the right roots

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf ☠ ~~~~2x(x+6)=1(x-5) }\) correct?

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf 2x^2+12x=x-5 }\) \(\Huge\color{blue}{ \sf 2x^2+11x+5=0 }\) solve this quadratic form me please

OpenStudy (anonymous):

(x + 5) (2x + 1)